40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7775     Entry time: Sun Dec 2 00:37:49 2012
Author: Den 
Type: Update 
Category: SUS 
Subject: TT cable problem 

 This week I've got all TT stuff baked and today was testing eddy current damping and electronics.

In the beginning everything was good: ring magnets fit mirror holder holes and their interaction with actuation magnets is strong enough to keep damping magnets in the wholes. I've put the frame horizontally and kicked it, magnets were still in the whole. Brackets also fit to the TT frame.  

DSC_4945.JPG     DSC_4946.JPG

I've tested eddy current dumping during ring down measurements, it was strong enough.

DSC_4947.JPG    DSC_4948.JPG

Then I started to test electronics. I've provided signal to TT1 channels and could see it in the clean room. But then things went terrible. I just could not connect TT cables to OSEMS, there is not enough space in the OSEM for the connector to plug in.

DSC_4949.JPG     DSC_4952.JPG

Connector should be machines to be more narrow. There is actually no reason for a connector to have this shape. I think it was designed to fit perfectly the OSEM frame but turned out to be ~0.5 mm wider then it should be.

ELOG V3.1.3-